# An alternative theory of magnetic flux tubes in strong fields via axion origin photons

### v2

@@ I Introduction
In this regard, let us return to the introduction of our section. Of all the known concepts that play a significant role in understanding the connection between the energy transfer of a magnetic flux tube and the darkness of sunspots, we will consider what we believe to be the most significant. It is based on the Parker-Biermann cooling effect (E. N. Parker, 1955a; L. Biermann, 1941; E. N. Parker, 1979a) in strong magnetic fields, which explains how the result of the suppression of Parker’s convective heat transfer manifests itself in the lower part of the magnetic flux tube (see Fig. 3).
@@
In order to understand the physics of the Parker’s suppression of the convective heat transport in strong magnetic fields, we need to tun to the dark matter axions, born in the core of the Sun. With solar axions and the existence of a magnetic O-loop inside the MFT near the tachocline, the answer becomes very simple. When a magnetic O-loop is formed inside the MFT near the tachocline through the Kolmogorov turbulent cascade (see Fig. 4), the high-energy photons from the radiation zone experience the axion-photon oscillations in this O-loop, and the so-called axions of photonic origin appear under the sunspot. This means that the cooling effect of Parker-Biermann exists due to the disappearance of barometric equilibrium (E. N. Parker, 1974a) and, as a consequence, the manifestation of the photon free path (Rosseland length; see Fig. B.3 in V. D. Rusov et al. (2021)) from the tachocline to the photosphere, which is confirmed by axions of photonic origin after photon-axion oscillations in the O-loop (see Fig. 3).
@@
Figure 3: Topological effects of magnetic reconnection inside the magnetic tubes with the “magnetic steps” (Fig. B.1 in V. D. Rusov et al. (2021)). The left panel shows the temperature and pressure change along the radius of the Sun from the tachocline to the photosphere (J. Bahcall & M. Pinsonneault, 1992), $L_{MS}$ is the height of the magnetic shear steps. At $R\sim 0.72~R_{Sun}$ the vertical magnetic field, which is developed from the horizontal part of the magnetic field (step) with the participation of the O-loop through the well-known Kolmogorov turbulent cascade (see Fig. 4), reaches $B_{z}\sim 3600$ T, and the magnetic pressure $p_{ext}=B^{2}/8\pi\simeq 5.21\cdot 10^{13}~erg/cm^{3}$ (J. Bahcall & M. Pinsonneault, 1992). The very cool regions along the entire convective zone caused by the Parker-Biermann cooling effect have the virtually zero internal gas pressure, i.e. the maximum magnetic pressure in the magnetic tubes. The narrow “purple” rings between the O-loop and the tube walls ($\rho_{ext}=\rho_{int}^{0}\gg\rho_{int}$) with the Parker-Bierman cooling effect inside, are a very important result of the existence of convective heating $(dQ/dt)_{2}$ in Section II.1.
@@
On the other hand, high-energy photon fluxes coming from the radiation zone through the thin “ring” part of the magnetic tube, i.e. between the tachocline and the “ring” (or more precisely between the magnetic wall and the O-loop of the tube; see Figs. 3 and 4), and which is applied without the Parker-Biermann cooling effect, allows to determine the convective heating $(dQ/dt)_{2}$ (see Sect. II.1).
@@
Figure 4: Kolmogorov turbulent cascade (A. Kolmogorov, 1941; A. N. Kolmogorov, 1968, 1991) and primary magnetic reconnection (which differs sharply from the secondary (see Fig. 5b) in the lower layers inside the unipolar magnetic tube (Fig. B.2 in V. D. Rusov et al. (2021)). Common to these various turbulent systems is the presence of the inertial region of Kolmogorov, through which the energy is cascaded from large to small scales. In this case, dissipative mechanisms (as a consequence of the “primary” magnetic reconnection) overcome the turbulent energy during plasma heating.
@@
As described by E. N. Parker (2009, 1994); H. C. Spruit et al. (1987); P. R. Wilson et al. (1990), the upward convection flow around the rising $\Omega$-loop brings its “legs” together in such a way that the magnetic field reconnection occurs across this loop. This cuts off the magnetic loop from the azimuthal magnetic field, turning it into an O-loop (see Fig. 4 in (E. N. Parker, 1994) and Fig. 3 in (H. C. Spruit et al., 1987)). After that the azimuthal magnetic field restores its initial configuration and becomes ready for another process with $\Omega$-loop.
Let us now make some important remarks on the turbulent reconnection, the $\Omega$-loop transformation into the O-loop by rapid “legs” closure, the restoration of the initial azimuthal field and the preconditions for another $\Omega$-loop formation in the same place. It is also necessary to explain the physical interpretation of the overshoot process near the tachocline and estimate the velocity $v_{rise}$ and time $\tau_{d}$ of the magnetic tube rise from the overshoot boundary layer – starting with the azimuthal magnetic flux strength of $B_{tacho}\sim 4\cdot 10^{7}~G$ (see Eq. A.1 in (V. D. Rusov et al., 2021)).
One of the final goals of this section is to determine the general regularities of the theory of magnetic flux tubes in strong fields of the order of $\sim 10^{7}~G$ using photons of axion origin – the so-called universal Ballegooijen-Fan-Fisher model (Sec. II.1), which are generated by the magnetic buoyancy of practically empty tubes, and as a consequence, their rise from the tachocline to the solar surface (Figs. 3-5). Another goal is the essence of the physics of the primary and secondary reconnection of magnetic flux tubes in the lower layers of the tachocline (Sect. II.2), which is directly related to the observed features of the cycle of both double sunspot maxima (see Sect. II.2.2) and the tilt angle of Joy’s law (see Sect. II.2.1). Moreover, both effects (see Sect. II.1-II.2) are caused by the existence (of dark matter?) – solar axions generated in the core of the Sun.
II.1 Radiative $(dQ/dt)_{1}$ sec-reconnectionand convective $(dQ/dt)_{2}$ heatings and the buoyant rise of magnetic flux tubes and the phenomenon of solar axions
@@
The assumption that the virtually empty magnetic tubes (Fig. 3) are neutrally buoyant ($\rho_{int}=\rho_{ext}$ (E. N. Parker, 1994)) implies that the temperature inside these tubes is lower than that of the ambient medium (Fig. 3 and Fig. 7a). This leads to the heat inflow, and consequently, the flux tube rises up (see E. N. Parker (1975) or Sect. 8.8 in (E. N. Parker, 1979b)). For a horizontal tube with a cross-section of radius $a$ the rise velocity follows from the Parker’s analysis (E. N. Parker (1975, 1979b), also Eq. (60) in (A. A. van Ballegooijen, 1982)):
@@
where $H_{p}=\Re T_{ext}/g=p_{ext}/g\rho_{ext}=0.08R_{Sun}$ (E. Böhm-Vitense, 1958; H. C. Spruit, 1977; A. S. Brun et al., 2011) is the pressure scale height at the tachocline, $T_{ext}$ and $p_{ext}$ are the external gas temperature and pressure, $\delta\equiv Y=\nabla_{e}-\nabla_{ad}=-c_{p}^{-1}dS/d\xi=-c_{p}^{-1}H_{p}dS/dz$ is the dimensionless entropy gradient (see A. A. van Ballegooijen (1982); R. Smolec & P. Moskalik (2008, 2010)), $\nabla_{e}\equiv d\ln T_{e}/d\ln p_{e}$ and $\nabla_{ad}\equiv(\partial\ln T_{e}/\partial\ln p_{e})_{s}$ are the local and adiabatic temperature gradients in external and internal plasma (H. Spruit, 1974; A. A. van Ballegooijen, 1982; J. Christensen-Dalsgaard et al., 1995), $s$ is the specific entropy, $c_{p}$ is the heat capacity at constant pressure, and $\tau_{d}$ is the rise time of the radiative heating of the magnetic flux tube:
@@
where for the fully ionized gas $c_{p}=2.5\Re=2.5p_{ext}/\rho_{ext}T_{ext}$ ($\Re$ is the the equation of state $p_{ext}=\rho_{ext}\Re T_{ext}=nk_{B}T_{ext}$ (E. Böhm-Vitense, 1958)), $T(z)$ and $\rho(z)$ are the mean temperature and density; $k_{e}$ is the radiative heat conductivity (see Eq. (35) in A. A. van Ballegooijen (1982)); $\ell_{ov}\approx 0.37H_{p}$ (A. A. van Ballegooijen, 1982; J. Christensen-Dalsgaard et al., 2011) is the thickness of the overshoot layer; the total radiative energy flux $F_{tot}=L/(4\pi r^{2})$ depends on the Sun luminosity $L$; $g$ is the gravitational acceleration.
Next we apply the condition of hydrostatic equilibrium, $dp/dz=\rho g$, when the adiabatic temperature gradient $(dT/dz)_{ad}=g/c_{p}$ may be used, and the neutral buoyancy of the flux tube in the overshoot zone $(|\delta T|/T_{ext})^{-1}\sim\beta\equiv 8\pi p_{ext}/B^{2}$. This way we are able to estimate the time of the radiative and/or convective diffusion $\tau_{d}$ (see Eq. (4)) of the flux tube:
@@
Here $\nabla_{rad}=(\partial\ln T_{ext}/\partial\ln p_{ext})_{rad}$ is the radiative equilibrium temperature gradient; $(dQ/dt)_{1}$ is the rate of radiative heating, which only depends on the thermodynamic parameters $k_{e}$ and $T_{ext}$ of the ambient plasma, depending only on the radial distance from the Sun center (A. A. van Ballegooijen, 1982; J. Christensen-Dalsgaard et al., 1995).
As a result, it is not difficult to show that the van Ballegooijen model combining equations (3)-(7) gives the final expression for the rise time by radiation heating from the boundary layer of the overshoot to the solar surface,
@@
$\tau_{d}\approx\left.5\frac{1}{\beta}p_{ext}\middle/\left(\frac{dQ}{dt}\right)_{1}\right.\approx 0.84\cdot 10^{8}~sec\approx 2.5~years,$ (9)
by means of $1/\beta=B^{2}/8\pi\rho_{ext}\approx 2.9\cdot 10^{-5}$, which is identical to the magnetic field $B\approx B_{eq}\sim 2\cdot 10^{5}~G$ (see $B^{2}/8\pi=p_{ext}=nk_{B}T_{ext}\approx 1.665\cdot 10^{13}~erg\cdot cm^{-3}$ (at $0.8R_{Sun}$ (see e.g. Fig. 3)) and $(dQ/dt)_{1}\approx 29.7~erg\cdot cm^{-3}\cdot s^{-1}$ (see Eq. (18) in Y. Fan & G. Fisher (1996)).
@@
$v_{rise}=H_{p}\nabla_{ad}\frac{1}{p_{ext}}\left(\frac{dQ}{dt}\right)_{1}\left(|\delta|+0.12\frac{B^{2}}{8\pi p_{ext}}\right)^{-1},~~~\nabla_{ad}=\nabla_{e}\simeq 0.4,$ (10)
@@
by means of $H_{p}\approx 5.6\cdot 10^{4}~km$ and $(|\delta|+0.12B^{2}/8\pi p_{ext})=(3/5)\times B^{2}/8\pi p_{ext}+0.12\times B^{2}/8\pi p_{ext}\approx 0.72\times 2.9\cdot 10^{-5}$ (see Eqs. (58-60) in (A. A. van Ballegooijen, 1982)), are almost identical to the equations (29)-(30) of Y. Fan & G. Fisher (1996).
@@
Surprisingly, these equations (Eqs. (8) and (10)) are almost identical to equations (29) and (30) from Y. Fan & G. Fisher (1996), which arise from completely different equations (60)-(61) of J. Bahcall & M. Pinsonneault (1992).
@@
Here the second term $(dQ/dt)_{2}$ represents a convective diffusion across the flux tube that is due to the temperature difference ($\delta T\equiv T-T_{e}$) between the flux tube and the external plasma (see (Y. Fan & G. Fisher, 1996)).
@@
$v_{rise}=H_{p}\nabla_{ad}\frac{1}{p_{ext}}\frac{dQ}{dt}\left(-\delta+0.12\frac{B^{2}}{8\pi p_{ext}}\right)^{-1},~~~\nabla_{ad}=\nabla_{e}\simeq 0.4,$ (16)
is the general case of the so-called universal model of the van Ballegooijen-Fan-Fisher, which is completely identical to equation (27) of the model by Y. Fan & G. Fisher (1996):
$\frac{d\Delta\rho}{dt}=\rho_{e}\frac{v_{rise}}{H_{p}}\left[\delta+\left(\frac{1}{\gamma}-\frac{2}{\gamma^{2}}\right)\frac{1}{\beta}-\frac{1}{\gamma}\frac{\Delta\rho}{\rho_{e}}\right]+\frac{2}{\gamma\beta}\rho_{e}\left[\frac{\partial(v\cdot I)}{\partial s}-\vec{v}\cdot\vec{k}\right]+\frac{\rho_{e}}{p_{e}}\nabla_{ad}\left[\left(\frac{dQ}{dt}\right)_{1}+\left(\frac{dQ}{dt}\right)_{2}\right],$ (17)
where in case of quasi-static rising, it is necessary apply ($d\Delta\rho/dt$) and $\Delta\rho$ to zero (see equation (27) of Y. Fan & G. Fisher (1996)). Furthermore, according to Y. Fan & G. Fisher (1996), for simplicity, the lifting of a uniform horizontal magnetic flux tube in the region of a plane-parallel overshoot region is considered, under which the conditions $\partial(v\cdot I)/\partial s=0$ and $\vec{v}\cdot\vec{k}=0$ can be used in equation (27) of Y. Fan & G. Fisher (1996): $I\equiv\partial r/\partial s$ is the unit vector tangential to the flux tube, and $k\equiv\partial^{2}r/\partial s^{2}$ is the tube’s curvature vector.
It should also be recalled that, taking into account the presence of subadiabaticity $\delta<0$ (see Eq. (58) in (A. A. van Ballegooijen, 1982), Eq. (27) in (Y. Fan & G. Fisher, 1996), and also Fig. 6) of the overshoot tachocline and using, according to A. A. van Ballegooijen (1982), non-local mixing length theory (H. C. Spruit & A. A. van Ballegooijen, 1982; E. Böhm-Vitense, 1958; S. Z. Ali & S. Dey, 2020; M. A. Weber & Y. Fan, 2015), it can be shown that thin, neutrally buoyant flow tubes, stable in a stratified medium, provided that its field strength $B$ is smaller than a critical value $B_{c}$ (A. A. van Ballegooijen, 1982), which is approximately given by
@@
where we used the MFT heating rate which consists of the radiative heating rate $(dQ/dt)_{1}$ and convective heating rate $(dQ/dt)_{2}$ (see Eqs. (II.1)–(19), Eqs. (13) and (19) in (Y. Fan & G. Fisher, 1996) and Eqs. (8)–(9) in (M. A. Weber & Y. Fan, 2015)); the approximate ratio $|\delta T|/T_{ext}\sim 1/\beta$; the parameter $\alpha_{1}^{2}\approx 5.76$ (Y. Fan & G. Fisher, 1996; M. A. Weber & Y. Fan, 2015); $F_{rad}$ is the radiative energy flux (see Eq. (15) in (Y. Fan & G. Fisher, 1996), and also Eq. (6) and Fig. 1 in (M. A. Weber & M. K. Browning, 2016)); $\nabla_{e}\sim 1.287\nabla_{rad}$ (see Table 2 in (H. C. Spruit, 1977)); $H_{p}/a$ is the factor for the lower convection zone (see (R. M. Roettenbacher et al., 2016)), where it was previously believed that $a=(\Phi/\pi B_{tacho})^{1/2}\leqslant 0.1H_{p}$ (see 3.2.1 in (R. M. Roettenbacher et al., 2016)) with an average value of typical magnetic flux of $\Phi\sim 10^{20}-10^{22}~Mx$ (Y. Fan et al., 1993).
From this, we understand that it is with the strong fields that a sharp increase in convective heating $(dQ/dt)_{2}$ is necessary in the tube, which simultaneously leads to a sharp decrease in the outer width of the ring $a\equiv a_{conv}$ (between the O-loop and the tube walls in Fig. 6) in a practically empty magnetic flux tube (see Fig. 3 and Fig. 6). In this case, the area of the “ring” of the magnetic tube is approximately equal to $2\pi r\times a_{conv}$, where $r$ is the radius of the magnetic tube, and $a_{conv}$ is the width of the “ring”.

### v3

@@ II Physics of the theory of practically empty flux tubes in strong fields of the order $\sim 10^{7}~G$
Of all the known concepts that play a significant role in understanding the connection between the energy transfer of a magnetic flux tube and the darkness of sunspots, we will consider what we believe to be the most significant. It is based on the Parker-Biermann cooling effect [6, 7, 8] in strong magnetic fields, which explains how the result of the suppression of Parker’s convective heat transfer manifests itself in the lower part of the magnetic flux tube (see Fig. 2).
@@
To understand the physics of the Parker’s suppression of the convective heat transfer in strong magnetic fields, we need to tun to the dark matter axions originating from the Sun’s core. Taking into account solar axions and the existence of a magnetic O-loop within the magnetic field near the tachocline (see Fig. 3d), the answer becomes very simple. When a magnetic O-loop forms within the magnetic field near the tachocline via a turbulent Kolmogorov cascade (see Fig. 3), dark matter axions and high-energy photons from the radiation zone to the tachocline undergo both axion-photon and photon-axion oscillations in this O-loop during the magnetic field in the tachocline, during which so-called of axion origin photons and photon origin axions emerge under a sunspot. This means that the Parker-Birman cooling effect exists due to the disappearance of barometric equilibrium [10, 11] and, as a consequence, the manifestation of the photon mean free path (Rosseland length; see Fig. B.3 in [14]) from the tachocline to the photosphere, which is confirmed by the existence of both axion-photon oscillations during the formation of axion origin photons, and simultaneously by the existence of photon-axion oscillations during the formation of photon origin axions during MTF in the tachocline (see Fig. 2).
@@
Figure 2: Topological effects of magnetic reconnection inside the magnetic tubes with the “magnetic steps” (Fig. B.1 in [14]). The left panel shows the temperature and pressure change along the radius of the Sun from the tachocline to the photosphere [49], $L_{MS}$ is the height of the magnetic shear steps. At $R\sim 0.72~R_{Sun}$ the vertical magnetic field, which is developed from the horizontal part of the magnetic field (step) with the participation of the O-loop through the well-known Kolmogorov turbulent cascade (see Fig. 3), reaches $B_{z}\sim 3600$ T, and the magnetic pressure $p_{ext}=B^{2}/8\pi\simeq 5.21\cdot 10^{13}~erg/cm^{3}$ [49]. The very cool regions along the entire convective zone caused by the Parker-Biermann cooling effect have the virtually zero internal gas pressure, i.e. the maximum magnetic pressure in the magnetic tubes. The narrow “purple” rings between the O-loop and the tube walls ($\rho_{ext}=\rho_{int}^{0}\gg\rho_{int}$) with the Parker-Bierman cooling effect inside, are a very important result of the existence of convective heating $(dQ/dt)_{2}$ in Section II.1.
@@
On the other hand, high-energy photon fluxes coming from the radiation zone through the thin “ring” part of the magnetic tube, i.e. between the tachocline and the “ring” (or more precisely between the magnetic wall and the O-loop of the tube; see Figs. 2 and 3), and which is applied without the Parker-Biermann cooling effect, allows to determine the convective heating $(dQ/dt)_{2}$ (see Sect. II.1).
@@
Figure 3: Kolmogorov turbulent cascade [50, 51] and primary magnetic reconnection (which differs sharply from the secondary (see Fig. 4) in the lower layers inside the unipolar magnetic tube (Fig. B.2 in [14]). Common to these various turbulent systems is the presence of the Kolmogorov’s inertial region, through which the energy is cascaded from large to small scales. In this case, dissipative mechanisms (as a consequence of the “primary” magnetic reconnection) overcome the turbulent energy during plasma heating.
@@
As described by [52, 53, 54], the upward convection flow around the rising $\Omega$-loop brings its “legs” together in such a way that the magnetic field reconnection occurs across this loop. This cuts off the magnetic loop from the azimuthal magnetic field, turning it into an O-loop (see Fig. 4 in [52] and Fig. 3 in [53]). After that the azimuthal magnetic field restores its initial configuration and becomes ready for another process with $\Omega$-loop.
Let us now make some important remarks on the turbulent reconnection, the $\Omega$-loop transformation into the O-loop by rapid “legs” closure, the restoration of the initial azimuthal field and the preconditions for another $\Omega$-loop formation in the same place. It is also necessary to explain the physical interpretation of the overshoot process near the tachocline and estimate the velocity $v_{rise}$ and time $\tau_{d}$ of the magnetic tube rise from the overshoot boundary layer – starting with the azimuthal magnetic flux strength of $B_{tacho}\sim 4\cdot 10^{7}~G$ (see Eq. A.1 in [14]).
One of the final goals of this section is to determine the general regularities of the theory of magnetic flux tubes in strong fields of the order of $\sim 10^{7}~G$ using photons of axion origin – the so-called universal Ballegooijen-Fan-Fisher model (Sec. II.1), which are generated by the magnetic buoyancy of practically empty tubes, and as a consequence, their rise from the tachocline to the solar surface (Figs. 2-4). Another goal is the essence of the physics of the primary and secondary reconnection of magnetic flux tubes in the lower layers of the tachocline (Sect. II.2), which is directly related to the observed features of the cycle of both double sunspot maxima (see Sect. II.2.2) and the tilt angle of Joy’s law (see Sect. II.2.1). Moreover, both effects (see Sect. II.1-II.2) are caused by the existence (of dark matter?) – solar axions generated in the core of the Sun.
II.1 Radiative $(dQ/dt)_{1}$ and convective $(dQ/dt)_{2}$ heatings with the concern of the Parker-Biermann cooling effect and axions of photon origin
@@
The assumption that the virtually empty magnetic tubes (Fig. 2) are neutrally buoyant ($\rho_{int}=\rho_{ext}$ [52]) implies that the temperature inside these tubes is lower than that of the ambient medium (Fig. 2 and Fig. 4a). This leads to the heat inflow, and consequently, the flux tube rises up (see [55] or Sect. 8.8 in [56]). For a horizontal tube with a cross-section of radius $a$ the rise velocity follows from the Parker’s analysis ([55, 56], also Eq. (60) in [57]):
@@
where $H_{p}=\Re T_{ext}/g=p_{ext}/g\rho_{ext}=0.08R_{Sun}$ [58, 59, 60] is the pressure scale height at the tachocline, $T_{ext}$ and $p_{ext}$ are the external gas temperature and pressure, $\delta\equiv Y=\nabla_{e}-\nabla_{ad}=-c_{p}^{-1}dS/d\xi=-c_{p}^{-1}H_{p}dS/dz$ is the dimensionless entropy gradient (see [57]), $\nabla_{e}\equiv d\ln T_{e}/d\ln p_{e}$ and $\nabla_{ad}\equiv(\partial\ln T_{e}/\partial\ln p_{e})_{s}$ are the local and adiabatic temperature gradients in external and internal plasma [61, 62, 63], $s$ is the specific entropy, $c_{p}$ is the heat capacity at constant pressure, and $\tau_{d}$ is the rise time of the radiative heating of the magnetic flux tube:
@@
where for the fully ionized gas $c_{p}=2.5\Re=2.5p_{ext}/\rho_{ext}T_{ext}$ ($\Re$ is the the equation of state $p_{ext}=\rho_{ext}\Re T_{ext}=nk_{B}T_{ext}$ [58]), $T(z)$ and $\rho(z)$ are the mean temperature and density; $k_{e}$ is the radiative heat conductivity (see Eq. (35) in [57]); $\ell_{ov}\approx 0.37H_{p}$ [62, 63] is the thickness of the overshoot layer; the total radiative energy flux $F_{tot}=L/(4\pi r^{2})$ depends on the Sun luminosity $L$; $g$ is the gravitational acceleration.
Next we apply the condition of hydrostatic equilibrium, $dp/dz=-\rho g$, when the adiabatic temperature gradient $(dT/dz)_{ad}=g/c_{p}$ may be used, and the neutral buoyancy of the flux tube in the overshoot zone $(|\delta T|/T_{ext})^{-1}\sim\beta\equiv 8\pi p_{ext}/B^{2}$. This way we are able to estimate the time of the radiative and/or convective diffusion $\tau_{d}$ (see Eq. (4)) of the flux tube:
@@
Here $\nabla_{rad}=(\partial\ln T_{ext}/\partial\ln p_{ext})_{rad}$ is the radiative equilibrium temperature gradient; $(dQ/dt)_{1}$ is the rate of radiative heating, which only depends on the thermodynamic parameters $k_{e}$ and $T_{ext}$ of the ambient plasma, depending only on the radial distance from the Sun center [57, 62].
As a result, it is not difficult to show that the van Ballegooijen model combining equations (3)-(7) gives the final expression for the rise time by radiation heating $(dQ/dt)_{1}$ from the boundary layer of the overshoot to the solar surface,
@@
$\tau_{d}\approx\left.5\frac{1}{\beta}p_{ext}\middle/\left(\frac{dQ}{dt}\right)_{1}\right.\approx 0.81\cdot 10^{8}~sec\approx 2~years,$ (9)
by means of $1/\beta=B^{2}/8\pi\rho_{ext}\approx 2.9\cdot 10^{-5}$, which is identical to the magnetic field $B\approx B_{eq}\sim 2\cdot 10^{5}~G$ (see $B^{2}/8\pi=p_{ext}=nk_{B}T_{ext}\approx 1.665\cdot 10^{13}~erg\cdot cm^{-3}$ (at $0.8R_{Sun}$ (see e.g. Fig. 2)) and $(dQ/dt)_{1}\approx 29.7~erg\cdot cm^{-3}\cdot s^{-1}$ (see Eq. (18) in [64]).
@@
$v_{rise}=H_{p}\nabla_{ad}\frac{1}{p_{ext}}\left(\frac{dQ}{dt}\right)_{1}\left(|\delta|+0.12\frac{B^{2}}{8\pi p_{ext}}\right)^{-1},$
$\nabla_{ad}=\nabla_{e}\simeq 0.4,$ (10)
@@
by means of $H_{p}\approx 5.6\cdot 10^{4}~km$ and $(|\delta|+0.12B^{2}/8\pi p_{ext})=(3/5)\times B^{2}/8\pi p_{ext}+0.12\times B^{2}/8\pi p_{ext}\approx 0.72\times 2.9\cdot 10^{-5}$ (see Eqs. (58-60) in [57]), are almost identical to the equations (29)-(30) of [64].
@@
Surprisingly, these equations (Eqs. (8) and (10)) are almost identical to equations (29) and (30) from [64], which arise from completely different equations (60)-(61) of [49].
@@
Here the second term $(dQ/dt)_{2}$ represents a convective diffusion across the flux tube that is due to the temperature difference ($\delta T\equiv T-T_{e}$) between the flux tube and the external plasma (see [64]).
@@
$v_{rise}=H_{p}\nabla_{ad}\frac{1}{p_{ext}}\frac{dQ}{dt}\left(-\delta+0.12\frac{B^{2}}{8\pi p_{ext}}\right)^{-1},$
$\nabla_{ad}=\nabla_{e}\simeq 0.4,$ (16)
is the general case of the so-called universal model of the van Ballegooijen-Fan-Fisher, which is completely identical to equation (27) of the model [64]:
$\frac{d\Delta\rho}{dt}$ $=\rho_{e}\frac{v_{rise}}{H_{p}}\left[\delta+\left(\frac{1}{\gamma}-\frac{2}{\gamma^{2}}\right)\frac{1}{\beta}-\frac{1}{\gamma}\frac{\Delta\rho}{\rho_{e}}\right]+$ (17)
$\frac{2}{\gamma\beta}\rho_{e}\left[\frac{\partial(v\cdot I)}{\partial s}-\vec{v}\cdot\vec{k}\right]+\frac{\rho_{e}}{p_{e}}\nabla_{ad}\left[\left(\frac{dQ}{dt}\right)_{1}+\left(\frac{dQ}{dt}\right)_{2}\right],$
where in case of quasi-static rising, it is necessary apply ($d\Delta\rho/dt$) and $\Delta\rho$ to zero (see equation (27) of [64]). Furthermore, according to [64], for simplicity, the lifting of a uniform horizontal magnetic flux tube in the region of a plane-parallel overshoot region is considered, under which the conditions $\partial(v\cdot I)/\partial s=0$ and $\vec{v}\cdot\vec{k}=0$ can be used in equation (27) of [64]: $I\equiv\partial r/\partial s$ is the unit vector tangential to the flux tube, and $k\equiv\partial^{2}r/\partial s^{2}$ is the tube’s curvature vector.
It should also be recalled that, taking into account the presence of subadiabaticity $\delta<0$ (see Eq. (58) in [57], Eq. (27) in [64], and also Fig. 5) of the overshoot tachocline and using, according to [57], non-local mixing length theory [57, 65, 58, 66, 67], it can be shown that thin, neutrally buoyant flow tubes, stable in a stratified medium, provided that its field strength $B$ is smaller than a critical value $B_{c}$ [57], which is approximately given by
@@
where we used the MFT heating rate which consists of the radiative heating rate $(dQ/dt)_{1}$ and convective heating rate $(dQ/dt)_{2}$ (see Eqs. (19)–(21), Eqs. (13) and (19) in [64] and Eqs. (8)–(9) in [67]); the approximate ratio $|\delta T|/T_{ext}\sim 1/\beta$; the parameter $\alpha_{1}^{2}\approx 5.76$ [64, 67]; $F_{rad}$ is the radiative energy flux (see Eq. (15) in [64], and also Eq. (6) and Fig. 1 in [68]); $\nabla_{e}\sim 1.287\nabla_{rad}$ (see Table 2 in [56]); $H_{p}/a$ is the factor for the lower convection zone (see [69]), where it was previously believed that $a=(\Phi/\pi B_{tacho})^{1/2}\leqslant 0.1H_{p}$ (see 3.2.1 in [57], 5.6.1 in [69]) with an average value of typical magnetic flux of $\Phi\sim 10^{20}-10^{22}~Mx$ [70, 71, 72].
From this, we understand that it is with the strong fields that a sharp increase in convective heating $(dQ/dt)_{2}$ is necessary in the tube, which simultaneously leads to a sharp decrease in the outer width of the ring $a\equiv a_{conv}$ (between the O-loop and the tube walls in Fig. 4a) in a practically empty magnetic flux tube (see Fig. 2 and Fig. 4a). In this case, the area of the “ring” of the magnetic tube is approximately equal to $2\pi r\times a_{conv}$, where $r$ is the radius of the magnetic tube, and $a_{conv}$ is the width of the “ring”.